# NoteLLM-2: Multimodal Large Representation Models for Recommendation

### v1

@@ Appendix G Performance of Zero-shot Multimodal Representation
Appendix H Case Study
Figure 6: The visualization cases of some methods. Each line is a case, which the left two columns are the query note and target note, the other columns are the top-$1$ ranked notes for each method. We also show the target note ranking.
We show some cases for multimodal I2I note recommendation as shown in Figure 6. In the first case, both the query note and the target note discuss a guide to Oahu, an island in Hawaii. MTomato-Base + NoteLLM-2 successfully ranks the target note in the top position. While METER Co-attn and MTomato-Base also recall relevant notes in the first place, they differ slightly. One talks about an unpleasant experience in Oahu, and the other discusses an overall trip plan for Hawaii. The second example illustrates a note pair about the Pagani car model. Although all baselines recall notes about various car models, these notes discuss many different types of cars. MTomato-Base + NoteLLM-2 ranks the Pagani model car in the first place, which can be attributed to both the image and the text content. The third case discusses how to make good coffee and coffee tricks. LLM-based methods recall relevant notes about the tools needed to make coffee and how to do it. In contrast, METER Co-attn, which has a weaker understanding of text, only recommends coffee beans. The last case involves a short note pair. The query note is about an AI drawing for China’s National Day, and the target note is about a mooncake drawn by AI, which relates to the Chinese traditional festival, the Mid-Autumn Festival. We find that MTomato-Base + NoteLLM-2 can understand the query note image about Chinese culture and recommends notes talking about Chang’e, a figure from Chinese folklore associated with the Mid-Autumn Festival. Tomato and MTomato-Base only focus on the AI painting, rather than the image content.

### v2

@@ Appendix D Performance of Zero-shot Multimodal Representation
Appendix E Model Performance on Long Query and Target Pairs
To comprehensively evaluate our method’s effectiveness, we conduct tests on Long Query and Long Target Pairs. Long notes are defined as those exceeding $165$ tokens (about $10\%$ of test notes). The test dataset contains 2,228 long query pairs and 2,177 long target pairs. The results are shown in Table 11. Due to the excellent long-context understanding of LLMs, the performance on long pairs is significantly higher than on overall pairs, as long pairs provide more information. Additionally, our method improves upon the basic method for long pairs, which shows that enhancing the focus on visual information is also important for long pairs.
Table 11. Performance of different methods based on three MLRMs on Long Query Pair and Long Target Pair (%).
  Base Model Method Long Query Pair Long Target Pair
R@100 R@1k R@10k R@100 R@1k R@10k
  MTomato-Base basic method 80.41 95.06 98.93 78.21 94.07 98.56
mICL 80.70 95.50 99.37 79.64 94.12 98.81
late fusion 80.85 95.79 99.41 79.84 94.36 99.01
NoteLLM-2 81.57 95.55 99.32 79.34 95.05 99.25
only late fusion 80.80 95.30 99.32 80.23 94.76 98.91
Omni-Retrieval 79.98 94.68 98.98 77.77 93.13 98.67
  MQwen-Base basic method 80.80 95.02 99.23 79.99 94.47 98.86
mICL 82.01 95.31 99.37 80.98 94.37 99.01
late fusion 82.16 95.50 99.42 81.07 94.51 99.11
NoteLLM-2 82.01 95.99 99.47 81.47 95.01 99.31
only late fusion 81.96 95.93 99.46 82.11 94.56 99.06
Omni-Retrieval 80.95 95.55 99.23 79.25 94.32 98.72
  MQwen-bigG basic method 82.78 96.17 99.60 82.31 95.75 99.01
mICL 83.36 96.17 99.61 82.51 95.80 99.16
late fusion 82.65 95.93 99.56 82.26 95.16 99.21
NoteLLM-2 82.74 96.32 99.56 81.82 95.80 99.01
only late fusion 82.64 95.79 99.46 80.78 95.35 99.25
Omni-Retrieval 82.30 96.03 99.41 80.63 95.16 99.11